OpenStudy (kainui):

You can order a pizza with up to 8 possible toppings. Which is greater? The number of possible pizzas you can get with only 3 toppings or the number of possible pizzas you can get with 5 toppings?

OpenStudy (vishweshshrimali5):

1st case = \(\large{^8C_3 \times 3!}\) 2nd case = \(\large{^8C_5\times 5!}\) Since, \(\large{^nC_r = ^nC_{n-r}}\) Thus: \[\large{^8C_3 = ^8C_5}\] Thus: 2nd case > 1st case

OpenStudy (vishweshshrimali5):

I have assumed these things: 1. All toppings are different. 2. When you say number of toppings on pizza, you mean the `exact` number of toppings.

OpenStudy (vishweshshrimali5):

3. Complete pizza = pizza base + topping(s) (with order of toppings important)

OpenStudy (vishweshshrimali5):

4. Different order of toppings mean different complete pizza

OpenStudy (anonymous):

Why would the order of the toppings be important? :3

OpenStudy (vishweshshrimali5):

I know - "many assumptions" but that's what I can think of now.

OpenStudy (vishweshshrimali5):

Well I have not eaten many pizzas, so I am leaving that to you all

OpenStudy (vishweshshrimali5):

If order is not important then both have the same number

OpenStudy (anonymous):

By the way, vishweshshrimali5 there's an easier way of writing \[\Large ^nC_r\cdot r!\] And that's \[\Large ^nP_r\] ^.^

OpenStudy (vishweshshrimali5):

Yeah I realised that while typing that stuff but left it as it gave the same answers :) But thanks.

OpenStudy (kainui):

I had not considered that the order in which the toppings are placed matters. So yes, you are right haha.

OpenStudy (anonymous):

Would it matter, though? Does the pizza look like some sort edible archery target? haha

OpenStudy (kainui):

It was just an amusing "trick" question I came across and thought it was fun but all the smart people came and knew I couldn't slide it past you guys haha.

OpenStudy (vishweshshrimali5):

It worked (I guess) :D !! So, .... Kainui.... when are you going to offer me and Peter the pizzas we prepared

OpenStudy (anonymous):

I'd love a slice :3

OpenStudy (vishweshshrimali5):

*number of pizzas we both found

OpenStudy (kainui):

brb lol

OpenStudy (vishweshshrimali5):

Hey thats not a pizza :P

OpenStudy (kainui):

who wants the ones with 5 toppings and who wants the ones with 3 toppings?

OpenStudy (anonymous):

I want the one pizza with all 8 toppings.

OpenStudy (vishweshshrimali5):

Same here ^^^^^^^

OpenStudy (kainui):

How about I give you 1 pizza with 0 toppings, it's pretty much the same.

OpenStudy (anonymous):

:/

OpenStudy (kainui):

lol